OpenStudy (anonymous):

Choose the reaction for which the heat contribution to the change in internal energy is negative AND the work contribution to the change in internal energy is positive

OpenStudy (anonymous):

\[ 4NH3(g) + 5O2(g) \rightarrow 4NO(g) + 6H2O(g) ΔrH° = -908 kJ mol–1\] \[2H2S(g) + SO2(g) \rightarrow3S(s) + 2H2O(g) ΔrH° = -729 kJ mol–1\] \[ CH4(g) + H2O(g) \rightarrow CO(g) + 3H2(g) ΔrH° = +206 kJ mol–1\] \[ N2(g) + O2(g) \rightarrow 2NO(g) ΔrH° = +90 kJ mol–1\]